Yah, right Not in your wildest dreams

A sample article for testing the adlart package^{∗}

Your Name^{†}  Paul Lato^{‡}

11 May 1965 Last updated: 05/02/2009

Abstract

This is the abstract.

Keywords. sample, article

AMS Subject Classifications. 00a00

Short Title. Sample article

1. A section with a math symbol: (P, ω)

This is a section. Look at the centred header and bold math mode.

1.1. A subsection. This is a subsection. Notice the header is inline with the text.

A subsubsection. This is a subsubsection. The header is inline with the text,

and there is no number.  Here's a marginal note.

2. Another section

Okay, now I need to babble a bit so we can get onto another page and you can see the running header. I suppose I could have just done a \newpage or something, but this is ever so much more entertaining.

I should point out that if you do not like the formatting I use, and I would not blame you if you do not, then you are welcome to change it. However, in the manner of the GNU public license, you are not allowed to keep the same name of the style file. But I intend for these macros to be more a guide to assist you in doing certain things you may find helpful.

∗This title should change for your article!

†Your title, Your address

Email: your.username@your.machine, URL: http://your.machine/~your.username/

You should fill in your own details here!

‡Generally Wise Fellow, The Academy, Athens, Greece

Current Address: Resting Place for Dead Philosophers

Email: follower@socrates.org, URL: http://academy.org/~plato/

Research supported in part by GPF grant number PHIL-1.

Okay, I think this is enough to get the running header on page 2. Notice that the author names as they appear in the running header are the optional argument of the \author command. Similarly, if I were able to keep babbling for another page, on page 3 the running header would be the optional argument

of \title.

Enjoy!

Here’s a bibliography. Notice it is \small.

References

[1] Michel Goossens, Frank Mittelbach, and Alexander Samarin. The L^{A}TEX Companion. Addison Wesley, Addison Wesley, 1994.